OpenStudy (anonymous):

which of the following is a solution to the equation X^2-6X+11=0 a. 3-square root(2i) b.the equation has no solutions real or conplex c.3+2square root(2i) d.3-2i e.3-2 square root(10) f.none of the others

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)